# AnchorDP3: 3D Affordance Guided Sparse Diffusion Policy for Robotic Manipulation

### v1

@@ 1 Introduction
As part of the RoboTwin Dual-Arm Collaboration Challenge [^1] https://robotwin-benchmark.github.io/, which centers on learning dual-arm visuomotor policies from simulation alone, participants are tasked with developing a single end-to-end model that can follow language instructions and operate under extreme scene randomization. The RoboTwin pipeline automatically constructs diverse simulation environments by generating 3D assets from RGB images using AIGC techniques, assigning affordance coordinates via 3D semantic models, and synthesizing expert trajectories with large language models. Figure 1 illustrates examples of diverse 3D simulation scenes generated by the RoboTwin system, covering a range of manipulation tasks. Despite these advantages, our participation in the RoboTwin Dual-Arm Collaboration Challenge revealed several key limitations that hinder the development of scalable, generalizable end-to-end policies. In particular, we identified three core challenges: perceptual ambiguity in cluttered scenes, task interference under shared representation learning, and inefficient action modeling for long-horizon, multi-step tasks.
@@ 3.2 Observation processing
    where $\boldsymbol{\mu}_{k}=\frac{1}{|\mathcal{N}_{k}|}\sum_{\mathbf{p}_{j}\in\mathcal{N}_{k}}\mathbf{p}_{j}$. The normal $\mathbf{n}_{k}$ corresponds to the eigenvector of $\mathbf{C}_{k}$ with smallest eigenvalue.
@@
Down sampling. To handle computational constraints, we apply Farthest Point Sampling (FPS) to select $N=4096$ points:
@@
Figure 3: Simulator-Supervised Semantic Segmentation
@@
This automated process generates precise point-level annotations without manual labeling, leveraging the simulator’s complete scene knowledge. Next, we train a lightweight PointNet++ model to predict the critical object labels with 11-dimensional downsampled point cloud $\mathcal{S}$ as inputs and per-point binary prediction $\hat{l}_{k}\in[0,1]$ as outputs.
@@
Multi-task training protocol. All encoders and the action expert are trained end-to-end using mixed-task batches containing trajectories from all 8 categories. For each trajectory, only the encoder corresponding to its task receives gradient updates. The shared diffusion action expert $A$ processes the feature vector $\mathbf{f}_{k}$ from the encoder and updates its parameters. Task-specific encoder weights remain isolated with no parameter sharing This approach enables efficient knowledge transfer to the shared action model while preventing negative interference between distinct manipulation strategies.
Architectural simplicity justification. Consistent with findings in 3D Diffusion Policy (DP3) literature, we observed that complex point cloud networks (e.g., PointNet++, PointNeXt) significantly hindered training stability. The minimalist MLP architecture provided best convergence behavior compared to hierarchical point processors. This design choice prioritizes robust gradient flow and computational efficiency—critical factors given the 24GB GPU memory constraint.
@@ 3.3 Affordance-Anchored Keypose Diffusion
Crucially, these poses are geometrically anchored to object affordances through linear transformations, e.g., PG = affordance centroid + pre-grasp offset. Therefore, the keyposes are also denoted as action anchors. A grasp-place trajectory can be represented as PG-TGO-TGC-PG-PP-TPC-TPO-PP-HM in chronological order. For other robotic manipulation skills, they can also be divided according to similar principles. The reformation of action space creates a compressed action space where the policy only needs to learn 10-30 sparse action anchors per task rather than thousands of dense actions. The neural networks do not explicitly learn the names and tags of each anction anchor, but implicitly learn to generate a list of multiple future action anchors through end-to-end diffusion training.
@@
Full-state supervised training. We use a simple and unified L2 loss for all the 8$\times$32 output elements, where 32 is the augmented action dimension and 8 is the predict horizon. Although the only useful output is located at the top left corner 1$\times$14 because only the first action anchor is executed without referring to end poses, the full state supervision for all the elements is beneficial for converging.
@@ 3.4 Large-scale dataset construction based on action anchors
Conventional dataset construction for robotic policies captures dense state-action pairs at full control frequency (typically 20-25Hz). We observed this approach incurs significant redundancy, exhibiting kinematic monotonicity where actions are predictable through simple inertial extrapolation. Beyond computational waste, this redundancy: 1) Dilutes the representation of critical kinematic inflection points; 2) Limits environmental diversity due to rendering/storage bottlenecks; 3) Encourages policies to learn motion inertia rather than affordance-driven decisions. To overcome these limitations, we implement a resource-optimized sparse trajectory collection protocol, as shown in Figure 4. While the total data volume remains similar, our proposed method can significantly increase the number of trajectories while reducing data frames per trajectory, resulting in a larger number of individual trajectories.
Figure 4: RoboTwin dataset constructron based on action anchors
Two-phase data collection. Our data collection strategy for each trajectory is divided into two distinct phases: a rollout phase and a render phase. The rollout phase is executed solely on the CPU, running the expert policy without rendering. This phase serves a dual purpose: it rapidly filters out configurations that are impossible to complete the task, such as objects not being on the table, and, for successfully completed tasks, records the specific time steps where actions are designated as action anchors. Then the render phase constructs the sparse trajectory dataset using the proposed strategy.
Sparse trajectory dataset construction. Within a single trajectory, $\mathbf{O}_{t}$ represents the observation, including point cloud data and 32-dimensional proprioception, while $a_{t}$ denotes the control target under the current observation, also 32-dimensional. Instead of traditional sequence-based method that saves all $a_{t}$, our method only save the action anchors $a_{Ak}$. Each frame of observation is directly pointed to a list of future action anchors $[a_{A(k+1)},...,a_{A(k+n)}]$. Under this strategy, it can be observed that rendering every observation frames between two key action time points is inefficient because their labels are identical. Therefore, we employ a random sampling strategy, rendering only two randomly chosen frames between the two consecutive key action time points. By reducing the identical-labeled observations, the size of a trajectory is significantly reduced. The number of trajectories with diverse environments can be increased by 14.5 times within the same computational budget and dataset size. The increased exposure to different environment settings contributes to the improvement of the model’s accuracy.
@@ 4 Conclusion
This work addresses the critical challenges of dual-arm robotic manipulation in unstructured environments through AnchorDP3, a novel diffusion policy framework that achieved state-of-the-art performance in the RoboTwin Challenge. Our solution directly tackles three fundamental limitations identified during the competition. First, perceptual ambiguity in cluttered scenes is resolved through our simulator-supervised semantic segmentation module, which leverages rendered ground truth to inject affordance awareness into point cloud representations. Second, multi-task interference is mitigated via task-conditioned feature encoders, enabling specialized feature extraction while maintaining a unified action expert. Third, inefficient long-horizon action modeling is overcome by our affordance-anchored keypose diffusion approach, which predicts geometrically meaningful sparse keyposes supervised with full kinematic state. These innovations collectively enabled AnchorDP3 to achieve a 98.7% average success rate under extreme scene randomization - the highest result in the RoboTwin Challenge. By seamlessly integrating with procedurally generated simulation workflows, AnchorDP3 provides a compelling pathway toward scalable real2sim2real transfer for complex manipulation tasks. Future work will explore sim-to-real transfer and extension to dynamic environments.

### v2

@@ 1 Introduction
As part of the RoboTwin Dual-Arm Collaboration Challenge [^1] https://robotwin-platform.github.io/, which centers on learning dual-arm visuomotor policies from simulation alone, participants are tasked with developing a single end-to-end model that can follow language instructions and operate under extreme scene randomization. The RoboTwin pipeline automatically constructs diverse simulation environments by generating 3D assets from RGB images using AIGC techniques, assigning affordance coordinates via 3D semantic models, and synthesizing expert trajectories with large language models. Figure 1 illustrates examples of diverse 3D simulation scenes generated by the RoboTwin system, covering a range of manipulation tasks. Despite these advantages, our participation in the RoboTwin Dual-Arm Collaboration Challenge revealed several key limitations that hinder the development of scalable, generalizable end-to-end policies. In particular, we identified three core challenges: perceptual ambiguity in cluttered scenes, task interference under shared representation learning, and inefficient action modeling for long-horizon, multi-step tasks.
@@ 3.2 Observation processing
    where $\boldsymbol{\mu}_{k}=\frac{1}{|\mathcal{N}_{k}|}\sum_{\mathbf{p}_{j}\in\mathcal{N}_{k}}\mathbf{p}_{j}$. The normal $\mathbf{n}_{k}$ corresponds to the eigenvector of $\mathbf{C}_{k}$ with the smallest eigenvalue.
@@
Down sampling. To handle computational constraints, we apply Farthest Point Sampling (FPS) to select $N=4096$ points.
@@
Figure 3: Simulator-Supervised Semantic Segmentation.
@@
This automated process generates precise point-level annotations without manual labeling, leveraging the simulator’s complete scene knowledge. Next, we train a lightweight PointNet++ model to predict the critical object labels with the 11-dimensional downsampled point cloud $\mathcal{S}$ as inputs and per-point binary prediction $\hat{l}_{k}\in[0,1]$ as outputs.
@@
Multi-task training protocol. All encoders and the action expert are trained end-to-end using mixed-task batches containing trajectories from all 8 categories. For each trajectory, only the encoder corresponding to its task receives gradient updates. The shared diffusion action expert $A$ processes the feature vector $\mathbf{f}_{k}$ from the encoder and updates its parameters. Task-specific encoder weights remain isolated with no parameter sharing. This approach enables efficient knowledge transfer to the shared action model while preventing negative interference between distinct manipulation strategies.
Architectural simplicity justification. Consistent with findings in 3D Diffusion Policy (DP3) literature, we observed that complex point cloud networks (e.g., PointNet++, PointNeXt) significantly hindered training stability. The minimalist MLP architecture provided the best convergence behavior compared to hierarchical point processors. This design choice prioritizes robust gradient flow and computational efficiency-critical factors given the 24GB GPU memory constraint.
@@ 3.3 Affordance-Anchored Keypose Diffusion
Crucially, these poses are geometrically anchored to object affordances through linear transformations, e.g., PG = affordance centroid + pre-grasp offset. Therefore, the keyposes are also denoted as action anchors. A grasp-place trajectory can be represented as PG-TGO-TGC-PG-PP-TPC-TPO-PP-HM in chronological order. For other robotic manipulation skills, they can also be divided according to similar principles. The reformation of action space creates a compressed action space where the policy only needs to learn 10-30 sparse action anchors per task rather than thousands of dense actions. The neural networks do not explicitly learn the names and tags of each action anchor, but implicitly learn to generate a list of multiple future action anchors through end-to-end diffusion training.
@@
Full-state supervised training. We use a simple and unified L2 loss for all the 8$\times$32 output elements, where 32 is the augmented action dimension and 8 is the prediction horizon. Although the only useful output is located at the top left corner 1$\times$14 because only the first action anchor is executed without referring to end poses, the full state supervision for all the elements is beneficial for converging.
@@ 3.4 Large-scale dataset construction based on action anchors
Conventional dataset construction for robotic policies captures dense state-action pairs at full control frequency (typically 20-25Hz). We observed this approach incurs significant redundancy, exhibiting kinematic monotonicity where actions are predictable through simple inertial extrapolation. Beyond computational waste, this redundancy: 1) Dilutes the representation of critical kinematic inflection points; 2) Limits environmental diversity due to rendering/storage bottlenecks; 3) Encourages policies to learn motion inertia rather than affordance-driven decisions. To overcome these limitations, we implement a resource-optimized sparse trajectory collection protocol, as shown in Figure 4. While the total data volume remains similar, our proposed method can significantly reduce data frames per trajectory, resulting in a larger number of individual trajectories.
Figure 4: RoboTwin dataset construction based on action anchors.
Two-phase data collection. Our data collection strategy for each trajectory is divided into two distinct phases: a rollout phase and a render phase. The rollout phase is executed solely on the CPU, running the expert policy without rendering. This phase serves a dual purpose: it rapidly filters out configurations that are impossible to complete the task, such as objects not being on the table, and, for successfully completed tasks, records the specific time steps where actions are designated as action anchors. Then, the render phase constructs the sparse trajectory dataset using the proposed strategy.
Sparse trajectory dataset construction. Within a single trajectory, $\mathbf{O}_{t}$ represents the observation, including point cloud data and 32-dimensional proprioception, while $a_{t}$ denotes the control target under the current observation, also 32-dimensional. Instead of the traditional sequence-based method that saves all $a_{t}$, our method only saves the action anchors $a_{Ak}$. Each frame of observation is directly pointed to a list of future action anchors $[a_{A(k+1)},...,a_{A(k+n)}]$. Under this strategy, it can be observed that rendering every observation frame between two key action time points is inefficient because their labels are identical. Therefore, we employ a random sampling strategy, rendering only two randomly chosen frames between the two consecutive key action time points. By reducing the identical-labeled observations, the size of a trajectory is significantly reduced. The number of trajectories with diverse environments can be increased by 14.5 times within the same computational budget and dataset size. The increased exposure to different environmental settings contributes to the improvement of the model’s accuracy.
@@ 4 Conclusion
This work addresses the critical challenges of dual-arm robotic manipulation in unstructured environments through AnchorDP3, a novel diffusion policy framework that achieved state-of-the-art performance in the RoboTwin Challenge. Our solution directly tackles three fundamental limitations identified during the competition. First, perceptual ambiguity in cluttered scenes is resolved through our simulator-supervised semantic segmentation module, which leverages rendered ground truth to inject affordance awareness into point cloud representations. Second, multi-task interference is mitigated via task-conditioned feature encoders, enabling specialized feature extraction while maintaining a unified action expert. Third, inefficient long-horizon action modeling is overcome by our affordance-anchored keypose diffusion approach, which predicts geometrically meaningful sparse keyposes supervised with full kinematic state. These innovations collectively enabled AnchorDP3 to achieve a 98.7% average success rate under extreme scene randomization - the highest result in the RoboTwin Challenge. By seamlessly integrating with procedurally generated simulation workflows, AnchorDP3 provides a compelling pathway toward tackling complex manipulation tasks. Future work will explore sim-to-real transfer and extension to dynamic environments.